You have observed the following returns over time YearStock X Stock Y Market 2012 2013 2014 2015 2016 13% 1496 19 16 12% 10 12 20 15 Assume that the risk-free rate is 6% and the market risk premium is 5%. a. What are the betas of Stocks X and Y? Answer b. What are the required rates of return on Stocks X and Y? Answer C. What is the required rate of return on a portfolio consisting of 80% of Stock X and 20% of Stock Y? Answer
Answer #1

SEE IMAGE. ANY DOUBTS, FEEL FREE TO ASK. HAPPY TO HELP YOU. THUMBS UP PLEASE. THANK YOU

Home nert Page Layout Formulas Data Review View dd-Ins s Cut aCopy Σ AutoSum ー E ゴWrap Text в 1 프 . Ej-., Δ. : rーー 逻锂函MergeHome nert Page Layout Formulas Data Review View dd-Ins Cut Σ AutoSum ー E ゴWrap Text aCopy B 1 u. ,_a. ars-函Merge & Center,Home nert Page Layout Formulas Data Review View dd-Ins Cut E AutoSum ー E ゴWrap Text aCopy B า 프 . Ej-., Δ. : r_一 逻锂函Merge & Center. $, % , 弼,8 Paste Conditional Format CeInsert Delete Format Formatting as Table Styles2 Clear Sort &Find & Format Painter Clipboard Font Alignment Number Styles Cells Edting AA95 AB AC AD AE AF AG AH Al Al AK 80 81 82 83 84 85 86 87 ANSWER b REQUIRED RATE OF RETURN Rf +beta(Rm-Rf) 696 + 1.3471(596) FORX FOR X ke = ke = Rm-Rf-RISK PREMIUM-5% 12.74% (Rx) FOR Y FOR Y 6% + 0.6508(596) ke- ke = 9.25% (Ry) 89 90 91 92 93 94 95 96 97 98 ANSWER c PORTFOLIO RETURN wx 0.80 wy =0.20 PORTFOLIO RETURN PORTFOLIO RETURN wxRx + wy*Ay= 0.80*(12.74%) + 0.20*(9.25%) Wx Rx+Wy Ry 12.04% DEAR (Sheet2 | COV TIME SERIES corr REGRESSION CAMERAEXP RETURN MATRIK INTERVAL, NORMALHY 福 130% 06:13 05-01-2019